OpenStudy (anonymous):

Hello, What is the general solution of the following differential equation? (1 + (x^3/y)sin^2x)dx + (1/y)(x + (1/cos^2(2y))dy = 0; I used the integrating factor, but I'm stuck at the following step: u(x) [-x^3sin^2x + 1] = u'(x) [x + (1/cos^2(2y)] Thanks in advance!

OpenStudy (abb0t):

can you rewrite it using the equation editor on the bottom? It's hard to see what you have there, but it looks like an exact differential equation.

OpenStudy (anonymous):

yes, im trying to make it into one. \[u(x) [-x ^{3}\sin ^{2} x + 1] = u'(x) [x + (1/\cos ^{2}(2y)]\]

OpenStudy (abb0t):

No, I mean the original differential equation.

OpenStudy (anonymous):

The original is a non-exact equation, but I'm trying to use the integrating factor method to make it into an exact equation.

OpenStudy (anonymous):

Any idea on how to simplify what i have to solve for u(x)?

OpenStudy (abb0t):

Did you use substitution method to put it into a first order differential equation? I don't quite understand what you're doing. Sorry. lol. That's why I'm having troubles helping you right now.

OpenStudy (anonymous):

First, I checked if the equation was an exact one. I found out that it wasn't, so I multiplied both sides of the equation by u(x) = integrating factor. This gave me a new equation which lets us know that it is in fact an exact equation.

OpenStudy (anonymous):

Now i just have to solve for the integrating factor.

OpenStudy (anonymous):

The original equation is: \[(1 + (x^3/y)\sin^2x)dx + (1/y)(x + (1/\cos^2(2y))dy = 0\]

OpenStudy (anonymous):

\[ \left(1+{x^3\over y}\sin^2x\right)+{1\over y}\left(x+{1\over\cos^2(2y)}\right){dy\over dx}=0 \] is this what it is!!!!

OpenStudy (anonymous):

can read it properly...

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

now i can nvermind.

OpenStudy (anonymous):

it was just showing a bunch of symbols before.

OpenStudy (anonymous):

ok now?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so, what was your step for this?

OpenStudy (anonymous):

do you have any idea how to go on from where im stuck? or any other suggestions?

OpenStudy (anonymous):

u(x)[−x 3 sin 2 x+1]=u ′ (x)[x+(1/cos 2 (2y)]

OpenStudy (anonymous):

could you write that clearly like I did.

OpenStudy (anonymous):

i did before, let me try again

OpenStudy (anonymous):

\[u(x) [-x ^{3}\sin ^{2} x + 1] = u'(x) \left[x +{1\over\cos ^{2}(2y)}\right]\]

OpenStudy (anonymous):

yes that is it.

OpenStudy (anonymous):

im trying to solve for the integrating factor u(x).

OpenStudy (anonymous):

problem 1) in the first paranthesis... something smells bad

OpenStudy (anonymous):

how would you solve it?

OpenStudy (anonymous):

\[M(x,y)=1+{x^3\over y}\sin^2x\qquad N(x,y)={x\over y}+{1\over y\cos^2(2y)}\\ M_y=-{x^3\over y^2}\sin^2x\qquad\qquad\quad N_x={1\over y}\]

OpenStudy (anonymous):

I got those values, therefore indicating that the equation is not exact.

OpenStudy (anonymous):

\[ N_x-M_y={1\over y}\left(1+{x^3\over y}\sin^2x\right)={M\over y}\\ \therefore \ln[u(y)]=\int{1\over y}dy=\ln y\\ \boxed{u(y) = y} \]

OpenStudy (anonymous):

The next part is very tedious especially to put on this IM area.

OpenStudy (anonymous):

why did u subtract?

OpenStudy (anonymous):

well thank you and im going to absorb this as soon as possible!

OpenStudy (anonymous):

then we create our new DE \[ (y+x^3\sin^2x)+\left[x+{1\over\cos^2(2y)}\right]{dy\over dx}=0 \] let us check if this is exact \[M(x,y)=y+x^3\sin^2x\qquad N(x,y)=x+{1\over\cos^2(2y)}\\ M_y=1\qquad\qquad\qquad N_x=1\qquad{\rm=exact}\\ F(x,y)=\int M\,dx+g(y)\qquad F(x,y)=\int N\,dy+g(x) \]

OpenStudy (anonymous):

1)\[F=\int\left[x+\sec^2(2y)\right]dy+g(x)\\ \boxed{\large F(x,y)=xy+{\tan(2y)\over2}+g(x)} \]

OpenStudy (anonymous):

but \(F_x=M\) \[ y+g'(x)=y+x^3\sin^2x\\\implies g(x)=\int(x^2\sin^2x)dx \]

OpenStudy (anonymous):

correction, make it a cube

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

this is the only worse part.. find g(x) and complete your F(x,y)

OpenStudy (anonymous):

i couldnt get that far.

OpenStudy (anonymous):

ill try to get there though; thanks!

OpenStudy (anonymous):

did you follow the steps?

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

yup, i had a prob with understanding why subtracted the partial derivatives earlier.

OpenStudy (anonymous):

you*

OpenStudy (anonymous):

to find the missing factor u(y) to make the equation exact.

OpenStudy (anonymous):

i get that but y subtract them.erhaps im missing something.

OpenStudy (anonymous):

we have to see what difference would make it a factor of one of these N(x,y) or M(x,y)

OpenStudy (anonymous):

subtraction gives you the missing portion :)

OpenStudy (anonymous):

i really appreciate it! thanks a bunch!

OpenStudy (anonymous):

welcome another bunch

OpenStudy (anonymous):

how do u give a badge?

OpenStudy (anonymous):

you did! when you clicked the "best response"

OpenStudy (anonymous):

haha ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

peace out

OpenStudy (anonymous):

np peace!